During summer shutdown further modifications of BEP were done. Two modern pickups were installed. Two copper mirrors for synchrotron radiation based beam diagnostics were changed due to insufficient quality. The new SR outlet for future material science experiments was equiped with movable water cooled absorber. At the end of the K-500 tunnel the Faraday cup was installed for absolute beam charge measurements.
The first beam returned to the complex only in October. After beam scrubbing at BEP and beam tests at VEPP-2000 the superconducting final focus (FF) solenoids were cooled down to helium temperatures in November. The same was done with CMD-3 detector solenoid. Finally the nominal "Rund Beam" regime was established and the beam was captured at the energy of 510 MeV on 22 Nov 2016.
The routine work on machine tuning was started with orbit correction and lattice correction. The very delicate procedure of beam-based alignment of solenoids was carried out in order to minimize the sthrength of dipole correctors, which suffer with poor field quality being embedded into quads. First luminosity of 5·1030cm-2s-1 at 510 MeV was observed on 15 Dec 2016. It is still twice smaller than record values of 2013.